A SPECTROSCOPIC AND THEORETICAL STUDY OF WEAK INTRAMOLECULAR OH⋯ π INTERAC-TIONS IN ALLYL CARBINOL AND METHALLYL CARBINOL

\underline{SIDSEL D. SCHROEDER}, KASPER MACKEPRANG, HENRIK G. KJAERGAARD, Department of Chemistry, University of Copenhagen, Universitetsparken 5, DK-2100 Copenhagen Ø.

The weak intramolecular OH⋯ π interactions in allyl carbinol and methallyl carbinol have been studied using a combination of NIR spectroscopy and theory. The third OH-stretching overtone region of vapor phase allyl carbinol and methallyl carbinol have been recorded with intracavity laser photoacoustic spectroscopy to study the effect of an enhanced OH⋯ π interaction in methallyl carbinol arising from the electron donating methyl group. Local mode calculations were employed to assign the observed bands. The OH-stretching transition frequency of methallyl carbinol was observed to be red shifted relative to the OH-stretching transition frequency of allyl carbinol. A red shift of the transition frequency is in this context normally interpreted as a signature of hydrogen bonding. Whether the OH⋯ π interaction can be categorized as a hydrogen bond will be discussed in this talk.